ID |
Date |
Author |
Type |
Category |
Subject |
9961
|
Fri May 16 09:46:05 2014 |
Steve | Update | PSL | pointing monitoring |
Quote: |
Tonight I noticed that the drop in PMC transmission was ~1V, more than the usual of ~0.5V from the daily drift.
While re-aligning on the table, I noticed that the misalignment was not from either of the steering mirrors; i.e. I has to walk them both to get the alignment back. This implies that the misalignment is generated far upstream. Maybe the the laser itself is moving. We need some updates from Steve's laser misalignment tracker.
|
I'd like to replace the paper target with IOO -QPD_POS so we can log it. |
9960
|
Fri May 16 00:25:53 2014 |
rana | Update | PSL | PMC realign | Tonight I noticed that the drop in PMC transmission was ~1V, more than the usual of ~0.5V from the daily drift.
While re-aligning on the table, I noticed that the misalignment was not from either of the steering mirrors; i.e. I has to walk them both to get the alignment back. This implies that the misalignment is generated far upstream. Maybe the the laser itself is moving. We need some updates from Steve's laser misalignment tracker. |
9959
|
Thu May 15 16:46:35 2014 |
ericq | Update | LSC | Possible Path to AO path | It's taken a lot of trial and error, but I've found a path through MATLAB loops that seems like it may be stable at all points.
CAVEAT: This doesn't give any indication as to why we weren't able to turn up the AO gain more last night, as far as I can tell, so it's not all good news.
However, it's still ok to at least have a plan that works in simulation... 
Based on the location of the optical resonance peak in the CARM plant, we estimate our CARM offset to be 200pm. I haven't simulated TFs there exactly, but do have 100pm and 300pm TFs. This procedure works in MATLAB starting at either, though 100pm is a little nicer than 300. MATLAB data and code is attached in a zip.
The steps below correspond to the attached figures: Bode plots and step response of the Loop at each step.
0. [Not Plotted] DCTrans sensing, MCL actuation on CARM. FMs1,2,3,5,8; UGF = 120. (DARM not considered at all)
- AO path just turned on. Crossover with MCL path ~ 3.5kHz.
- AO gain increased. Crossover ~ 500 Hz. There are now multiple UGFs! Handling all of these in a stable manner is tricky.
- AO gain increased. Crossover = 150Hz. [No simulations with a higher crossover survived the next steps]
- Compensation filter applied to MCL path; 1 real Zero at 105Hz and a pole at 1k. From a TF point of view, this is sort of like switching to REFLDC, but the SNR at low frequencies is probably better in TR signals at this point.
- CARM offset reduced to 30pm. (This smoothens out the optical plant resonance.)
- Overall gain increased by factor of 3. There is now just one UGF at a few kHz, above the optical resonance. From here, gain can be further increased, boosts can go on, offset can go way down. In reality, we should switch to a single error signal once we're back to one UGF, and go from there.
 
#4 Seems like the most sticky part. While both sides of this look stable as far as I can tell. I feel that flipping from the red phase curve to the teal might not actually be ok, since they are on either side of the bad phase of 0 degrees. It isn't immediately evident to me how to easily model the transitions between steps, rather than just the stability of of each step in the steady state. |
Attachment 3: AOCarmLoop.zip
|
9958
|
Thu May 15 15:10:02 2014 |
Steve | Update | PEM | seismic activities | Our only seismometer is at the east end now.
Atm1, Ditch Day morning puzzle. The gardener came after the freshman did leave and cut the grass with the lawn mower.
Atm2, Yesterday afternoon the Aztecs containers moved out. |
Attachment 1: DichDay.png
|
|
Attachment 2: Ditchday8am.jpg
|
|
Attachment 3: AztecsGone.jpg
|
|
9957
|
Thu May 15 02:52:51 2014 |
Jenne | Update | LSC | Started engaging the AO path, not getting all the way yet |
In addition to a transparent legend, we need the corresponding CM crossover measurements from DTT to compare with the Q-Mist-Loop model results. The xover tells us when the AO gain is high enough so that they can be ramped up together.
Also, I wonder how much power fluctuation we get from the large ALS DIFF noise and if that demands we get the TR signals normalized by POPDC. |
9956
|
Thu May 15 02:32:01 2014 |
Jenne | Update | LSC | Started engaging the AO path, not getting all the way yet | I tried many times this evening to engage the AO path, with limited success.
Q's new scripts worked really well, and so I have some transfer functions! To take these measurements, in ...../scripts/general/netgpib, I am running ./TFSR785 TFSR785_CARMloop_May2014.yml , where the file name is the name of my parameter file. The data, and the saved pdfs, are in /users/jenne/PRFPMI/CARM_loop_measurements/2014May14/ . For these measurements, the SR785 is hooked up to the "A" set of excitation and test points on the CM board.
All of these traces were taken while the IFO was PRFPMI, with PRCL and MICH on REFL33, DARM on ALS diff, and CARM on InvSqrtTrans. carm_cm_up.sh is up to date, through the echo "REFL_I should now be zero" (~line 111 in the script). All you need to do is set the beatnotes, and then run the script. Follow instructions in the prompt (such as "press enter to confirm PRMI is locked").

Here are my notes for the various times:
23:01:44 - MC IN2 = 0dB, CARM gain = 5.0
23:13:45 - MC IN2 = 10dB, CARM gain = 5
23:26:10 - MC IN2 = 6dB, CARM gain = 10 (after Q suggested increasing overall gain, rather than just AO path)
00:13:07 - MC IN2 = 6dB?, CARM = 6ish? don't remember exactly.
00:45:00ish, Realigned IFO using IR with arms.
01:03:17 - MC In2 = 0dB, CARM gain = 5
01:07:42 - MC IN2 = 8dB, CARM gain = 6.295 (AO went up to 6dB, then +1dB steps to both simultaneously using ezcastep C1:LSC-CARM_GAIN 1dB C1:IOO-MC_AO_GAIN 1 )
01:08:57 - MC IN2 = 10dB, CARM gain = 7.92447
01:10:08 - MC IN2 = 12dB, CARM gain = 9.97631
lockloss when trying to add 1 more dB to both.
01:41:36 - MC IN2 = 12dB, CARM gain = 9.97
lockloss when just MC IN2 up by 1dB, left CARM gain alone.
Other notes:
The 60Hz noise in TRY is back. Since I thought I remembered someone suggesting that it was leakage light from the exit sign, Koji went in and wrapped the end table in foil, however the lines are still present. |
9955
|
Thu May 15 01:42:07 2014 |
rana | Update | CDS | /frames space cleared up, daqd stabilized | Script seems to be working now:
nodus:~>df -h | grep frames
fb:/frames 13T 12T 931G 93% /frames
|
9954
|
Wed May 14 17:36:32 2014 |
ericq | Update | CDS | New netgpib scripts for SR785 | I have redone the SPSR785 (spectrum measurement) and TFSR785 (TF measurement) commands in scripts/general/netgpibdata. This was mostly motivated by my frustration with typing out either a ton of command line arguments, or rooting around in the script itself; I'd rather just have a static file where I define the measurement, and can keep track of easier.
They currently take one argument: a parameter file where all the measurement details are specified. (i.e. IP address, frequencies, etc.) There are a few template files in the same directory that they use as default. (Such as TFSR785template.yml)
If you call the functions with the option '--template', it will copy a template file into your working directory for you to modify as you wish. "SPSR785 -h" gives you some information as well (currently minimal, but I'll be adding more)
In the parameter file, you can also ask for the data to be plotted (and saved as pdf) when the measurement is finished. In SPSR785, and soon TFSR785, you can specify a directory where the script will look for reference traces to plot along with the results, presuming they were taken with the same measurement parameters and have the same filename stem.
I've tested both on Pianosa, and they seem to work as expected.
Todo:
- Add support for modifying some parameters at the command line
- Extend to the Agilent analyzer
- Maybe the analyzer settings written to the output file should be verified by GPIB query, instead of writing out the intended settings. (I've never seen them go wrong, though)
- Make sure that the analyzer has PSD units off when taking a TF. (Thought I could use resetSR785 for this, but there's some funkiness happening with that script currently.)
- Possibly unify into one script that sees what kind of analyzer you're requesting, and then passes of to the device/measurement type specific script, so we don't have to remember many commands.
Comments, criticism, and requests are very welcome.
(P.S. all the random measurement files and plots that were in netgpibdata are now in netgpibdata/junk. I feel like this isn't really a good place to be keeping data. Old versions of the scripts I changed are in netgpibdata/oldScripts) |
9953
|
Wed May 14 14:39:31 2014 |
Jenne | Update | SUS | New buttons on IFO_ALIGN screen | It's starting to get a little crowded, but I modified the IFO_ALIGN screen to have new buttons to show the aligned / misaligned state of each optic. Koji made a good point, and I left the old restore script functional so that if the slider is moved significantly, we can always go back gently to the burt restored value. I have removed the old misalign function though, since we shouldn't ever be using that again.

|
9952
|
Wed May 14 10:04:06 2014 |
Steve | Summary | SUS | oplev laser and temperature | ETMX oplev qpd gain has to be increased.
|
Attachment 1: TempRules.png
|
|
9951
|
Wed May 14 02:37:58 2014 |
Jenne | Update | LSC | No big locking news | Tonight was mostly cleaning up some scripts, including the re-writing the restore and align scripts for the optics.
The new script is in the same folder as the old one (/opt/rtcds/caltech/c1/medm/MISC/ifoalign/NewAlignSoft.py), but is not yet called from the align screen. However, I'm using it in the carm up and down scripts, and it works nicely for the PRM. I need to check that the offset value is okay for all the other optics (i.e. are they getting misaligned enough?), but then I'll have the new script called from the screen. The new script, per Rana's suggestion, does not touch the bias sliders. Rather, it puts an offset in the pitch filter banks in the coil driver output matrix-of-filter-banks. Then the misalign routine turns the offset on, and the restore routine turns the offset off. This way we have a nice ramp time, but don't have to do the weird calculation of number of steps to take as is done in the current script. Also, the "save" functionality will be obsolete, since we're never touching the bias sliders except for actual alignment needs.
I'm not sure what changed, other than the HEPA being on lower, but the Xarm ALS was much better behaved tonight. I was able to hang out around arm powers of ~1 for as long as I wanted.
I didn't try to hand over to digital REFLDC, but I was trying a few times to engage the AO path. With the CM board set to Plus, I hear hooting when MC IN2 is about 4dB. With the CM board set to Minus, I didn't hear hooting, but I lost lock when I went from 13dB to 14dB.
Also, I put the cables for the SR785 back to the "A" set of test points and excitation, so that I could take a closed loop transfer function. However, I don't know where the latest working scripts to make a remote measurement are, so maybe we can take some loop measurements tomorrow.
The carm_cm_up script is good (for tonight) up to the prompt "Press enter to indicate that it is okay to turn on MC2 LSC FM8". There are "read"s every step of the way, so it goes nice and slow, but it'll do everything for you except any last tweaks of the PRM alignment after the PRMI is locked. |
9950
|
Tue May 13 22:55:57 2014 |
Jenne | Summary | SUS | ETMX oplev: cleanup | I believe that the Xend aux laser was turned off earlier today, for Steve's work swapping out the oplev. When I went down there, the red "off" LED was illuminated, and the LCD screen was showing something. I pushed the green "on" button, and I immediately got green.
Also, I saw that the 24Hz roll mode was very rung up on ETMX. I looked at the FM5 "bounce roll" filter, and it had some old values, 12Hz and 18Hz for the resonant gains. All other optics have the proper 16Hz and 24Hz frequencies. I copied the BS oplev bounce roll filter over to ETMX pit and yaw (both were wrong), and loaded them in. The mode is starting to ring down. |
9949
|
Tue May 13 17:45:21 2014 |
rana | Update | CDS | /frames space cleared up, daqd stabilized |
Late last night we were getting some problems with DAQD again. Turned out to be /frames getting full again.
I deleted a bunch of old frame files by hand around 3AM to be able to keep locking quickly and then also ran the wiper script (target/fb/wiper.pl).
controls@pianosa|fb> df -h; date
Filesystem Size Used Avail Use% Mounted on
/dev/sda1 440G 9.7G 408G 3% /
none 7.9G 288K 7.9G 1% /dev
none 7.9G 464K 7.9G 1% /dev/shm
none 7.9G 144K 7.9G 1% /var/run
none 7.9G 0 7.9G 0% /var/lock
none 7.9G 0 7.9G 0% /lib/init/rw
none 440G 9.7G 408G 3% /var/lib/ureadahead/debugfs
linux1:/home/cds 1.8T 1.4T 325G 82% /cvs/cds
linux1:/ligo 71G 18G 50G 27% /ligo
linux1:/home/cds/rtcds
1.8T 1.4T 325G 82% /opt/rtcds
fb:/frames 13T 12T 559G 96% /frames
linux1:/home/cds/caltech/users
1.8T 1.4T 325G 82% /users
Tue May 13 17:35:00 PDT 2014
Looking through the directories by hand it seems that the issue may be due to our FB MXstream instabilities. The wiper looks at the disk usage and tries to delete just enough files to keep us below 95% full for the next 24 hours. If, however, some of the channels are not being written because some front ends are not writing their DAQ channels to frames, then it will misestimate the disk size. In particular, if its currently writing small frames and then we restart the mxstream and the per frame file size goes back up to 80 MB, it can make the disk full.
For now, I have modified the wiper.pl script to try to stay below 93%. As you can see by the above output of 'df', it is already above 96% and it still has files to write until the next run of wiper.pl 7 hours from now at. at 6 AM.
IF we assume that its writing a 75MB file every 16 seconds, then it would write 405 GB of frames every day. There is 559 GB free right now so we are OK for now. With 405 GB of usage per day, we have a lookback of ~12TB/405GB ~ 29 days (ignoring the trend files). |
9948
|
Tue May 13 17:31:32 2014 |
Jenne | Summary | SUS | ETMX oplev laser replaced: New oplev gains set | I took loop measurements of ETMX pit and yaw, and set the upper UGF to be ~6Hz for both. This required a pitch gain of 25, and a yaw gain of 16.
The spectra look similar to what they were before Steve did the swap.


|
9947
|
Tue May 13 17:03:05 2014 |
Steve | Summary | SUS | ETMX oplev laser replaced |
Quote: |
Quote: |
For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.
To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.
|
I'm in the process of filling this table
OPLEV
SERVO
|
300 ^
2:0
|
BR |
ELP |
RLP |
BOOST |
RES |
GAIN
|
QPD
COUNTS
|
QPD
mW
|
QPD
beam
OD
|
HE/NE
output
mW
|
%
back
on QPD
|
|
|
|
|
|
|
|
|
|
|
|
|
|
ETMY PIT |
FM1 |
FM5 |
|
55 |
|
|
-30 |
8,200 |
0.2 |
|
3.3 |
|
YAW |
FM1 |
FM5 |
|
55 |
|
|
-28 |
|
|
|
|
|
ETMX PIT |
FM1 |
FM5 |
35 |
|
|
|
4.4 |
900 |
0.2 |
|
1.7 |
|
YAW |
FM1 |
FM5 |
35 |
|
|
|
2.1 |
1,750 |
0.33 |
|
2.8 |
|
ITMY PIT |
FM1 |
FM5 |
|
|
|
3.3 |
52 |
14,400 |
0.4 |
|
9.5 |
|
YAW |
FM1 |
FM5 |
|
|
|
3.3 |
-46 |
|
|
|
|
|
ITMX PIT |
FM1 |
FM5 |
50 |
|
|
3.3 |
30 |
7,400 |
0.17 |
|
2.8 |
|
YAW |
FM1 |
FM5 |
50 |
|
|
3.3 |
-20 |
|
|
|
|
|
BS PIT |
FM1 |
FM5 |
35 |
|
|
3.3 |
9 |
2,800 |
0.05 |
|
2.8 |
|
YAW |
FM1 |
FM5 |
35 |
|
|
3.3 |
-9 |
|
|
|
|
|
PRM PIT |
FM1 |
FM5 |
55 |
|
FM7 |
3.3 |
7 |
3200 |
0.06 |
|
2.8 |
|
YAW |
FM1 |
FM5 |
55 |
|
FM7 |
3.3 |
-5 |
|
|
|
|
|
SRM PIT |
FM1 |
|
|
|
|
|
-20 |
|
|
|
9.5 |
|
YAW |
FM1 |
|
|
|
|
|
20 |
|
|
|
|
|
I should replace ETMX He/Ne laser
|
|
9946
|
Tue May 13 13:27:58 2014 |
Steve | Summary | SUS | Optical Lever Servos setting table |
Quote: |
For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.
To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.
|
I'm in the process of filling this table
OPLEV
SERVO
|
300 ^
2:0
|
BR
16,24
Hz
|
ELP |
RLP |
BOOST |
RES |
GAIN
|
QPD
COUNTS
|
QPD
mW
|
QPD
beam
OD
|
HE/NE
output
mW
|
%
back
on QPD
|
|
|
|
|
|
|
|
|
|
|
|
|
|
ETMY PIT |
FM1 |
FM5 |
|
55 |
|
|
-30 |
8,200 |
0.2 |
|
3.3 |
|
YAW |
FM1 |
FM5 |
|
55 |
|
|
-28 |
|
|
|
|
|
ETMX PIT |
FM1 |
FM5 |
35 |
|
|
|
4.4 |
900 |
0.2 |
|
1.7 |
|
YAW |
FM1 |
FM5 |
35 |
|
|
|
2.1 |
|
|
|
|
|
ITMY PIT |
FM1 |
FM5 |
|
|
|
3.3 |
52 |
14,400 |
0.4 |
|
9.5 |
|
YAW |
FM1 |
FM5 |
|
|
|
3.3 |
-46 |
|
|
|
|
|
ITMX PIT |
FM1 |
FM5 |
50 |
|
|
3.3 |
30 |
7,400 |
0.17 |
|
2.8 |
|
YAW |
FM1 |
FM5 |
50 |
|
|
3.3 |
-20 |
|
|
|
|
|
BS PIT |
FM1 |
FM5 |
35 |
|
|
3.3 |
9 |
2,800 |
0.05 |
|
2.8 |
|
YAW |
FM1 |
FM5 |
35 |
|
|
3.3 |
-9 |
|
|
|
|
|
PRM PIT |
FM1 |
FM5 |
55 |
|
FM7 |
3.3 |
7 |
3200 |
0.06 |
|
2.8 |
|
YAW |
FM1 |
FM5 |
55 |
|
FM7 |
3.3 |
-5 |
|
|
|
|
|
SRM PIT |
FM1 |
|
|
|
|
|
-20 |
|
|
|
9.5 |
|
YAW |
FM1 |
|
|
|
|
|
20 |
|
|
|
|
|
I should replace ETMX He/Ne laser |
Attachment 1: OLsums.png
|
|
9945
|
Tue May 13 03:30:10 2014 |
Jenne | Update | LSC | Locking activities - no progress | [Jenne, Rana, EricQ]
We tried a few times to engage the AO path while holding CARM on sqrtTrans and DARM on ALS, but failed every time. Since we cannot stably hold lock at arm powers of 1, even though we were able to do so early last week, we think that we have a problem (obviously). One noticeable thing is that while held with ALS, the Xarm transmission fluctuates almost full power.
As we were seeing late last week, the Xarm IR transmission while held with ALS was fluctuating wildly, whether we were locked on individual arms or on CARM/DARM.
Tonight we took some out of loop spectra, with different HEPA settings, to see how that affected things. It looks like HEPA at the nominal ~20% is okay, but anything higher than that starts to affect the Xarm beatnote sensing, while it mostly leaves the Yarm beatnote sensing okay. Perhaps this is something that isn't tightened enough in the Xarm path, or something on a skinny floppy mount that needs to be more secure.
I am still a little confused though, why we don't see large power fluctuations in *both* arms while using ALS to control CARM/DARM. Why are we not seeing this Xarm noise being fed back into the Yarm, through either the ETM via DARM, or common stuff via CARM actuating on MC2.
Note that the change at high frequency is because I switched from using non-DQ channels to DQ channels, so that's not anything to pay attention to. The noise reduction we see is below about 20Hz.

Rana pointed out that our POPDC level was very small. We don't have screens for them, but the DC PDs have the same analog whitening as the RF PD signals do. I changed C1:LSC-POPDC_WhiteGain.VAL from 0 to 11. Now our POPDC while locked on sidebands is about 8,000 counts.
We also swapped the cables between the SR785 and the CM board around. Now channel 2 of the 785 goes to TP2A, channel 1 goes to TP2B, and the source goes to EXCB. This is so that we can break the AO loop with the disable switch just after the slow/fast split, and look at the transfer function before we close the loop. |
9944
|
Tue May 13 00:46:58 2014 |
rana | HowTo | PSL | PMC relocking | The PMC runs out of range sometimes due to the daily temperature swing. The voltage swings up after sunset and then starts to swing down before sunrise. So when you relock the PMC at the beginning of the locking night, the mnemonic from the PMC is:
Sun Go Low, Lock Me Voltage Low. |
9943
|
Mon May 12 22:52:59 2014 |
Jenne | Summary | SUS | Optical Lever Filters are all different | We need to go back and have a look at all of our optical lever control filters, and make sure they make sense.
In particular, we should have a look at the ITMs, since they have a huge amount of motion around 10Hz.
Notes: ETMX shouldn't have that lower notch. The bounce mode Qs should be lowered.

|
9942
|
Mon May 12 22:42:19 2014 |
rana | Summary | SUS | Optical Lever QPD Sum trends: they're almost all too weak | For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.
To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts. |
Attachment 1: OL-FB.png
|
|
Attachment 2: arms_140512.pdf
|
|
9941
|
Mon May 12 14:42:25 2014 |
steve | Update | endtable upgrade | optical table enclosure wall proposal and table at ETMY |
Quote: |
Sanwiched wall as shown: 1" clear acrylic, 2 layers of 0.004" thick "window tint", 1 layer of 0.007" thermashield and 0.125" yellow acrylic
Visibility: 70 %, Transmission of 1064 nm 2-3 % at 0-50 degrees incident, power density ~ 0.7 W/mm2
Max power 100 mW
|
More details about this east end " acrylic + " enclosure ( optical table cover ) can be found elog entry 6210, 7194 and 7106
Window tinted layer transmission plot is below.
We have a film which may meet your requirements and the values are shown below:
Wavelength (nm) Transmission Reflectance(front) Reflectance (back)
1060 .0772 .604 .759
1070 .0723 .615 .772
These values are taken from the LBNL Optics 6 program and if you have access to that program, the NFRC ID for the film is 202. If you do not have access to the program, I have a captured the graph which may be of some help. I apologize for the appearance of the graph but someone at LBNL decided it would look better with a dark gray background – the yellow is the transmission curve, the blue is the reflectance (front) and the green is the reflectance (back).

The film is referred to as “Hilite 70” and has a 72% visible light transmission. These results were obtained with the film mounted on 1/8” clear glass.
I
Saint-Gobain Performance Plastics
www.solargard.com
Please consider your environmental responsibility before printing this email.
|
|
|
|
Attachment 1: Hilite70.jpg
|
|
Attachment 2: ETMY-ISCT.jpg
|
|
Attachment 3: RoscoThermashield.pdf
|
|
9940
|
Mon May 12 10:42:01 2014 |
rana | Update | SUS | some Arm maintenance | I ran the ASS/ADS for the arms because the X-arm was way out. There was also some problem with its locking due to bad ramps in FM2. I copied over the filters from YARM and then adjusted some of the ramps and thresh trigs in the filter file until the transients in POX got smaller. Basically, you should not really be ramping on Integrators. Secondly, we should do some testing when adjusting the filter parameters.
I hooked up the 4395 to the MC servo board OUT2 so that we can monitor the error point when the PCDRIVE goes nuts. |
9939
|
Fri May 9 21:18:51 2014 |
Koji | Update | Green Locking | Reverted X green light power | It is actually very tricky to measure the green power at the output of the doubling crystal as the IR often leaks into the measurement.
I checked the green beam powers on the X/Y/PSL tables.
CONCLUSION: There is no green beam which exceeds 5mW anywhere in the 40m lab.
Note: The temperature of the doubling crystal at the X end was optimized to have maximum green power. It was 36.3degC and is now 36.7degC.
X END:
When the angles of the wave plates are optimized, we have 539mW input to the doubling crystal.
With the Xtal temperature of 36.7degC, where the green power is maximized, the power right after
the harmonic separator (H.S.) was 9.6mW.
Xtal temp 36.7degC ~~~
|
--539mW@IR-->{Xtal}-->/-->9.6mW-->{Mirror}-->4.69mW-->{Mirror}-->4.54mW-->{Faraday}
(H.S.)
If we believe these 4.69mW and 4.54mW are purely from the green, we have 4.8mW right after the H.S.
This corresponds to the conversion efficiency of 1.6%/W (cf. theretical number 2%/W)
By disabling the heating of the crystal, we can reduce the green light by factor of 60. But still the reading right after the H.S. was 5.3mW
Xtal temp 29.2degC ~~~
|
--539mW@IR-->
{Xtal}-->/-->5.3mW-->{Mirror}-->285uW-->{Mirror}-->74.3uW-->{Faraday}
(H.S.)
Naively taking the difference, the green beam right after the H.S. is 4.4mW.
In either cases, the green power right after the oven is slightly less than 5mW.
Y END:
When the angles of the wave plates are optimized, we have 287mW input to the doubling crystal.
With the Xtal temperature of 36.0degC, where the green power is maximized, the power right after
the harmonic separator (H.S.) was 0.86mW.
Xtal temp 36.0degC ~~~
|
--287mW@IR-->{Xtal}-->/-->0.86mW-->
(H.S.)
When the temperature was shifted to 39.2degC, the reading after the H.S. was 70uW. Therefore the contamination by the IR is small
in this setup and we can believe the above reading in 70uW accuracy. This 0.86mW corresponds to the conversion efficiency of 1.2%/W.
PSL
The incident IR is 80mW. We have 170uW after the H.S., which corresponds to the conversion efficiency of 2.6%/W. Maybe there is some IR contamination?
From the vacuum chamber total 1mW of green is derivered when both arms are locked and aligned.
Thus the total green power at the PSL table is less than 5mW. |
9938
|
Fri May 9 14:01:24 2014 |
Jenne | Update | LSC | CM board boost turn-on checkout | Note: I thought I was editing elog 9935, but somehow this became a whole new entry. Either way, all the info is in here.
As part of checking the common mode board before we get too carried away with using it, I looked at the time series of the AO servo output when I turned on various boosts, or changed gain values. As it turns out, basically anything that I did caused glitches. Oooops.
I plugged a function generator to the IN1 port of the CM board, with a freq of 400Hz, and a voltage of 10mVpp (which is the smallest value that it would allow). I plugged the BNC version of the servo output into a 300MHz 'scope.
First I looked at "boost" and "super boost", and then I looked at various steps of the AO gain slider. For all of the button presses that gave me glitches, I saved .png's of the 'scope screen (on a floppy, so I'll have to fetch the data tomorrow...).
Both enabling, and disabling the "Boost" button gave me glitches.
For "Super Boost", I saw glitches for all of the steps, 0->1, 1->2, 2->3.
For the AO path, I only started at 0dB, and only captured screenshots of glitches when I increased the gain, since presumably that's when we'll care the most during acquisition. I found that going down in gain caused glitches at every step! For increasing the gain, steps from an odd number of dBs to an even number consistently caused glitches. Steps from an even number to an odd number occasionally caused glitches, but they weren't very common. For the steps that did cause glitches, some were worse than others (7dB to 8dB, 15 dB to 16 dB, and 23 dB to 24 dB seemed the worst.)
After my work, I put all of the cables back, so that we should be ready to utilize the CM board for locking this evening.
For posterity, here are the notes that I took while I was working - I'll make them more coherent when I fold them in with my images tomorrow. The "first .png, next, etc." are because the 'scope numbers them in order as a default.
1st png = boost enable, then disable
2nd png = super boost, start at 0, then 1, then 2, then 3
3rd png = AO gain from 1 to 0
4th is AO gain from 0 to 1 (happens less often than 1->0, which is every time I get a glitch)
Next is AO gain 1->2, got 2 glitches!
3->2 glitch often, 2->3 much less often
next is 2->3
next png is 3->4, 2 glitches with weird dip
4->5, rare
next png is 5->6
6->7 is rare
next png is 7->8, which is nasty!!
8->9 is rare
png 9->10
10->11 is rare
png 11-> 12, 3 glitches
12->13 rare
png 13->14, 2 glitches
14->15, rare
png 15->16, kind of nasty
png 17->18, 2 glitches
png 19->20, 3 glitches
png 21->22, 2 glitches
png 23->24, kind of nasty
png 25->26, 2 glitches
png 27->28, 3 glitches, at least
png 29->30, 2 glitches
The screenshot of the Boost enable / disable I'll have to re-take. Apparently I instead caught a screenshot of the list of files on the floppy...ooops.
This is a shot of enabling the Super Boosts. At the beginning, it's at "0", so no superboosts (also, regular boost was off). Then, I switch to "1", and the trace gets a little fuzzy. Then I switch to "2", and it gets very fuzzy. Then I switch to "3", and a lot of the fuzz goes away. There's a glitch at each transition.

The following screenshots are all of various steps of the AO gain slider. For all of these, both the "boost" and "super boosts" were off. Each screenshot is a single gain step, even if there are several glitches captured.
First, 0dB to 1dB:

Next, 1dB to 2dB:

2dB to 3dB:

3dB to 4dB:

While increasing the gain, I didn't find any more steps from an even to an odd number where I got a glitch. They would glitch when I undid that step (decreased the gain), but over ~5 trials for each increase, I didn't ever catch a glitch. The odd to even steps still had glitches while increasing the gain though.
5dB to 6dB:

7dB to 8dB:

9dB to 10dB:

11dB to 12dB:

13dB to 14dB:

15dB to 16dB:

17dB to 18dB:

19dB to 20dB:

21dB to 22dB:

23dB to 24dB:

25dB to 26dB:

27dB to 28dB:

29dB to 30dB:

|
9937
|
Fri May 9 11:23:11 2014 |
steve | Update | Green Locking | decreased X green light power | Green light power decreased from 3 mW to 1 mW at the ETMX-ISCT shutter. More later. |
9936
|
Fri May 9 04:51:13 2014 |
ericq | Summary | LSC | REFL_DC handoff didn't work last night |
Quote: |
Last night after checking cabling and turning on ISS, we tried several times to handoff to REFL_DC but it didn't work at all.
Some issues:
- The ISS was injecting a lot of very low frequency power fluctuations because of bad AC coupling.
- The SR560 @ LSC rack was saturating a lot with the x10 gain that Jenne and Rana put in; we turned it back to G = 1.
- The ISS was also saturating a lot. We turned it off around 4 AM, but still no success.
- The ALS sequence for finding the Red Resonance takes too long (~2 minutes), so we're trying a faster scheme tonight.
|
Still no success tonight
- We took CARM OLTFs at various CARM offsets and could clearly see the peak in the optical TF (in once case ~2.5kHz), which gave us an indication of our offset (~200pm)
- REFLDC effectively sees the same plant TF as the transmission signals plus a zero at ~110 Hz, at all offsets under 1nm, from my simulations; this pushes up the optical resonance and causes a loop instability when we try to handoff.
- We need to make the CARM OLTF steeper to suppress this instability, but also to make a good crossover with the AO path, which otherwise has too similar of a slope around the UGF, as we saw with our one arm test.
- We're thinking of trying to turn the AO path on with REFLDC while keeping the arms on SQRTINV signals. This may be tricky, but if we can get the loop bandwidth above the optical peak, it'll be a lot easier to deal with, and transfer digital control to REFLDC as well.
|
9935
|
Fri May 9 04:09:39 2014 |
Jenne | Update | LSC | CM board boost turn-on checkout | As part of checking the common mode board before we get too carried away with using it, I looked at the time series of the AO servo output when I turned on various boosts, or changed gain values. As it turns out, basically anything that I did caused glitches. Oooops.
I plugged a function generator to the IN1 port of the CM board, with a freq of 400Hz, and a voltage of 10mVpp (which is the smallest value that it would allow). I plugged the BNC version of the servo output into a 300MHz 'scope.
First I looked at "boost" and "super boost", and then I looked at various steps of the AO gain slider. For all of the button presses that gave me glitches, I saved .png's of the 'scope screen (on a floppy, so I'll have to fetch the data tomorrow...).
Both enabling, and disabling the "Boost" button gave me glitches.
For "Super Boost", I saw glitches for all of the steps, 0->1, 1->2, 2->3.
For the AO path, I only started at 0dB, and only captured screenshots of glitches when I increased the gain, since presumably that's when we'll care the most during acquisition. I found that going down in gain caused glitches at every step! For increasing the gain, steps from an odd number of dBs to an even number consistently caused glitches. Steps from an even number to an odd number occasionally caused glitches, but they weren't very common. For the steps that did cause glitches, some were worse than others (7dB to 8dB, 15 dB to 16 dB, and 23 dB to 24 dB seemed the worst.)
After my work, I put all of the cables back, so that we should be ready to utilize the CM board for locking this evening.
For posterity, here are the notes that I took while I was working - I'll make them more coherent when I fold them in with my images tomorrow. The "first .png, next, etc." are because the 'scope numbers them in order as a default.
1st png = boost enable, then disable
2nd png = super boost, start at 0, then 1, then 2, then 3
3rd png = AO gain from 1 to 0
4th is AO gain from 0 to 1 (happens less often than 1->0, which is every time I get a glitch)
Next is AO gain 1->2, got 2 glitches!
3->2 glitch often, 2->3 much less often
next is 2->3
next png is 3->4, 2 glitches with weird dip
4->5, rare
next png is 5->6
6->7 is rare
next png is 7->8, which is nasty!!
8->9 is rare
png 9->10
10->11 is rare
png 11-> 12, 3 glitches
12->13 rare
png 13->14, 2 glitches
14->15, rare
png 15->16, kind of nasty
png 17->18, 2 glitches
png 19->20, 3 glitches
png 21->22, 2 glitches
png 23->24, kind of nasty
png 25->26, 2 glitches
png 27->28, 3 glitches, at least
png 29->30, 2 glitches
Somehow, the images got put into a whole new entry, even though I thought I was editing this one. Anyhow, please see elog 9938. |
9934
|
Fri May 9 01:36:28 2014 |
rana | Summary | SUS | Optical Lever QPD Sum trends: they're almost all too weak | We want there to be ~16000 cts of signal per quadrant on the optical levers. I think that most of the QPDs have been modified to have 100k transimpedance resistors.
From the attached 90 day trend, you can see that the ETMX, BS, PRM, and SRM are really low. We should figure out if we need to change the lasers or if the coating reflectivities are just low.
Steve, can you please measure the laser powers with a power meter and then reply to this entry?
Another possibility is that we are just picking a dim beam and a brighter one is available. |
Attachment 1: OLtrend.png
|
|
9933
|
Thu May 8 17:25:17 2014 |
steve | Update | LSC | TRY 60Hz noise hunt | I worked at the ETMY-ISCT this morning and late afternoon. I will continue the 60 Hz noise hunt tomorrow. |
9932
|
Thu May 8 17:00:56 2014 |
rana, Q | Summary | LSC | REFL_DC handoff didn't work last night | Last night after checking cabling and turning on ISS, we tried several times to handoff to REFL_DC but it didn't work at all.
Some issues:
- The ISS was injecting a lot of very low frequency power fluctuations because of bad AC coupling.
- The SR560 @ LSC rack was saturating a lot with the x10 gain that Jenne and Rana put in; we turned it back to G = 1.
- The ISS was also saturating a lot. We turned it off around 4 AM, but still no success.
- The ALS sequence for finding the Red Resonance takes too long (~2 minutes), so we're trying a faster scheme tonight.
|
9931
|
Thu May 8 15:55:43 2014 |
jamie | Update | CDS | python issues |
Quote: |
On pianosa: The ezca.Ezca class somehow initializes with its prefix set to "C1:", even though the docstring says the default is None. This makes existing scripts act wonky, because they're looking for channels like "C1:C1:FO-BLAH".
In ligo/apps/linux-x86_64, I ran ln -sfn cdsutils-old cdsutils to get the old version back for now, so I don't have to edit all of our up/down scripts.
Also, Chiara can't find the epics package when I try to load Ezca. It exists in '/usr/lib/pymodules/python2.6/epics/__init__.pyc' on pianosa, but there is no corresponding 2.7 folder on chiara.
|
I just pushed a fix to ezca to allow for having a truly empty prefix even if the IFO env var is set:
controls@pianosa:~ 0$ ipython
Python 2.6.5 (r265:79063, Feb 27 2014, 19:43:51)
Type "copyright", "credits" or "license" for more information.
IPython 0.10 -- An enhanced Interactive Python.
? -> Introduction and overview of IPython's features.
%quickref -> Quick reference.
help -> Python's own help system.
object? -> Details about 'object'. ?object also works, ?? prints more.
In [1]: import ezca
In [2]: ezca.Ezca()
Out[2]: Ezca(prefix='C1:')
In [3]: ezca.Ezca(ifo=None)
Out[3]: Ezca(prefix='')
In [4]: ezca.Ezca(ifo=None).read('C1:LSC-DARM_GAIN')
Out[4]: 0.0
This is in cdsutils r232, which I just installed at the 40m. I linked it in as well, so it's now the default version. You will have to make a modification to any python scripts utilizing the Ezca object, but now it's a much smaller change (just in the invocation line):
-ca = ezca.Ezca()
+ca = ezca.Ezca(ifo=None)
|
9930
|
Thu May 8 14:37:02 2014 |
Jenne | Update | ASC | POP ASC QPD power | I was thinking about POP today, and wanted to know if there was something to be done to allow us to use the PRCL ASC for at least a little bit farther into arm power buildup.
Anyhow, I checked, and while PRMI is locked on sidebands (ETMs misaligned), POP DC is about 80 counts, and the power measured by the Ophir power meter is 24 microWatts.
We were on the 3rd gain setting for the QPD's power amplifier. I turned it down to the "2" option. (When at 4, the front panel light indicates saturation).
It's not clear to me what the gain settings mean exactly. I think that "1" means 4*10^3 V/A, and "6" means 4*10^6 V/A (On-Trak OT301 info site), but I don't know for sure how the gain changes for the settings 2-5. Anyhow, I have changed the digital gain for the ASC to be -0.063 from -0.023 for both pitch and yaw. |
9929
|
Thu May 8 02:03:51 2014 |
rana | Update | ISS | ISS: fuse was blown, repaired, loop back on | Back in November, Nic and Evan turned on an SR560 based ISS. It uses the PMC TRANS PD as the error signal and makes an AC coupled loop with 2 SR560's and then it drives the RF amplifier which drives the AOM upstream of the PMC.
This was the saturating SR560 under the PSL table that Steve found this week*. Tonight I found that the +24 V rack fuse for this was blown. I replaced the previous 2A fuse with a new 2A fuse (turned off the +/24 V Sorensens during this operation). I think all of the servo settings are basically the same as before, except that I'm using a gain of 10000 instead of 50000 on the first SR560. It was saturating otherwise. My guess is that the fuse blew many months ago and no one has noticed...
I checked the out of loop performance in MC_TRANS and in the IFO REFL_DC and there's some high frequency improvement with the loops on.
The main improvement, however, was in lowering the HEPA fan speed. This should only be turned up to Hurricane when you are working on the table. Similarly, those of us trying to lock at night, can't really trust that the HEPA is set to its nominal low setting of 20%. The whole difference in the MC_TRANS from 5-50 Hz is from this however, so we can use this ISS reference .xml as a way to see if the HEPA is up too high.
If we want to do better for RIN from 100-1000 Hz for improving the REFL_DC/CARM noise, we would have to think of how to improve the PMC_TRANS PD RIN.
* Steve gets +1 point for finding this, but then -3 points for not elogging. |
Attachment 1: ISS.pdf
|
|
9928
|
Thu May 8 01:33:21 2014 |
ericq | Update | CDS | python issues | On pianosa: The ezca.Ezca class somehow initializes with its prefix set to "C1:", even though the docstring says the default is None. This makes existing scripts act wonky, because they're looking for channels like "C1:C1:FO-BLAH".
In ligo/apps/linux-x86_64, I ran ln -sfn cdsutils-old cdsutils to get the old version back for now, so I don't have to edit all of our up/down scripts.
Also, Chiara can't find the epics package when I try to load Ezca. It exists in '/usr/lib/pymodules/python2.6/epics/__init__.pyc' on pianosa, but there is no corresponding 2.7 folder on chiara.
|
9927
|
Thu May 8 00:40:39 2014 |
ericq | Update | LSC | BNC vs. 2pin LEMO for AO | I've checked that the 2pin lemo connector that was run some time ago from the LSC rack to the MC board does indeed transmit signals. To try and evaluate its suitability I did the following:
- Generated a 5mVpp 1.3kHz signal with an SR785 and fed that into CM board In1, all boosts off, 0dB AO gain.
- Both BNC and LEMO connected to CM servo out
- One of BNC or LEMO connected to IN2 of MC servo, input gain of 30dB but disabled, OUT2 switched to AO and fed to Agilent spectrum analyzer.
- Terminated MC IN2 for comparison.
No real difference was seen between the two cases. The signal peak was the same height, width. 60Hz and harmonics were of the same amplitude. Here are the spectra out to 200k, they are very similar.

Mode cleaner was locked during this whole thing. This may interfere with the measurement, but is similar to the use case for the AO path. If ground loop / spurious noise issues keep occurring, it will be worthwhile to examine the noise of the CM and MC servo paths, inputs and outputs more carefully. |
9926
|
Wed May 7 23:30:21 2014 |
jamie | Update | CDS | cdsutils should be working now | Should be fixed now. There were python2.6 compatibility issues, which only show up on these old distros (e.g. ubuntu 10.04).
controls@pianosa:~ 0$ cdsutils read C1:LSC-DARM_GAIN
0.0
controls@pianosa:~ 0$ cdsutils --version
cdsutils 230
controls@pianosa:~ 0$
|
9925
|
Wed May 7 23:09:06 2014 |
rana, jamie | Summary | Computer Scripts / Programs | Ottavia back on network | After Jamie fixed the third party repo issue with Ottavia, he was able to upgrade it to Ubuntu 12.04 Precise Pangolin. But its network stopped working.
I tried to fix its issues by ifconfig and GUI, but what it really wanted was for me to put the network cable back into its eth0 slot. The eth1 network card appears not be working anymore.
All seems fine now. Next I will mount the shared user disk from linux1 and put in a .bashrc. |
9924
|
Wed May 7 22:47:33 2014 |
rana | Update | CDS | cdsutils updated to version 226: not working on pianosa or rossa | controls@rossa:~ 0$ cdsutils read C1:LSC-DARM_GAIN
Traceback (most recent call last):
File "/usr/lib/python2.6/runpy.py", line 122, in _run_module_as_main
"__main__", fname, loader, pkg_name)
File "/usr/lib/python2.6/runpy.py", line 34, in _run_code
exec code in run_globals
File "/ligo/apps/cdsutils-226/lib/python2.6/site-packages/cdsutils/__main__.py", line 57, in <module>
imp.load_module('__main__', f, pathname, description)
File "/ligo/apps/cdsutils-226/lib/python2.6/site-packages/cdsutils/read.py", line 32, in <module>
print ezca.Ezca(prefix).read(rest, as_string=args.as_string)
File "/ligo/apps/linux-x86_64/cdsutils-226/lib/python2.6/site-packages/ezca/cached.py", line 17, in __call__
key = (args, tuple(kwargs.viewitems()))
AttributeError: 'dict' object has no attribute 'viewitems'
|
9923
|
Wed May 7 17:10:59 2014 |
rana | Update | CDS | cdsutils updated to version 226 | This upgrade from Jamie has given us the new apps (avg, servo, and trigservo). We should figure out if there's a way to integrate Masayuki's work, so that we can have a 'cdsutils demod' function too. |
9922
|
Wed May 7 16:31:12 2014 |
jamie | Update | CDS | cdsutils updated to version 226 |
controls@pianosa:~ 0$ cd /opt/rtcds/cdsutils/trunk/
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ svn update
...
At revision 226.
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ make
echo "__version__ = '226'" >lib/cdsutils/_version.py
echo "__version__ = '226'" >lib/ezca/_version.py
...
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ make ligo-install
python ./setup.py install --prefix=/ligo/apps/linux-x86_64/cdsutils-226
...
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ ln -sfn cdsutils-226 /ligo/apps/linux-x86_64/cdsutils
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ exit
...
controls@pianosa:~ 0$ cdsutils --version
cdsutils 226
controls@pianosa:~ 0$
|
9921
|
Wed May 7 14:01:36 2014 |
steve | Update | LSC | TRY 60Hz noise hunt |
Quote: |
This is an effort to get rid of our ground loops by isolating the electronic components from the optical table.
Aluminum mounting base plates of Thorlabs BA2 and Newport B-2 were replaced by plates or post made out of delrin material.
This is an insulator. DELRIN base plates were installed 6 places. The oplev-qpd has Nylon base plate.
The NPRO and HE/NE lasers are not isolated from the table. S8 and S9
I'm not sure about the doubling oven S10 
The optical table is grounded at G11 through ~1 Mohms to the ETMY chamber.
Alignment touch up needed at all D-marked component!
|
Attachment appendix:
D: component delrin isolated
N: component nylon isolated ( or Delrin )
S: component shell is shorting to optical table (except oven)
G: optical table ground
I failed to maximize TRY the pds. |
9920
|
Wed May 7 04:01:44 2014 |
rana, jenne | Update | LSC | PRFPMI: Common Mode servo using REFL_DC ON, CARM offset still non-zero |
- With REFL_DC coupled into the CM board through an SR560 (with an offset subtractor), we were able to transition to use it as the CARM error signal.
- We reduced the CARM offset until the arm powers went up to ~13.
- We had the AO path turned on and the MCL/AO crossover was ~150 Hz.
- We saw the double cavity pole come in from HF down to ~1-2 kHz. The lock stayed stable like this.
- We've set the IMC overall gain higher by +4dB in the mcup script. That's -4 dB from Eric's max gain earlier today.
- We have some scripts now for this scripts/PRFPMI/ : camr_cm_down.sh and carm_cm_up.sh
- The sequence was ALS -> SqrtInv while digital with CARM -> MC2. Then we digital transition to REFL_DC using the CM board switch to put REFL_DC into the REFL11_I socket.
- REFL_DC is noisy, so we upped the SR560 gain by 10 and compensated.
Also, we found the PRM OL off and turned it back on. The ETMY was swinging a lot after lock loss, so we set its SUSPOS damping gain to match the ETMX and it stopped swinging so much.
Next up: more of the same, make this sequence more stable, turn on CARM OSC and watch the LOCKI outputs while we slowly ramp between signals.
Also, what should be the sign of the CARM offset ??? |
9919
|
Tue May 6 19:38:13 2014 |
Jenne | Update | LSC | Set up for PRFPMI CM locking | To get ready for the PRFPMI CM trials tonight, I put AS55's cables back to their nominal state, and now have REFL11 I going to IN1 of the CM board. The OUT1 of the CM board goes to the REFL11I whitening channel.
REFLDC was not disconnected in the last few days, so it is still set up for IN2 of the CM board, with an external offset adjust. |
9918
|
Tue May 6 18:32:14 2014 |
steve | Update | LSC | TRY 60Hz noise hunt | This is an effort to get rid of our ground loops by isolating the electronic components from the optical table.
Aluminum mounting base plates of Thorlabs BA2 and Newport B-2 were replaced by plates or post made out of delrin material.
This is an insulator. DELRIN base plates were installed 6 places. The oplev-qpd has Nylon base plate.
The NPRO and HE/NE lasers are not isolated from the table. S8 and S9
I'm not sure about the doubling oven S10 
The optical table is grounded at G11 through ~1 Mohms to the ETMY chamber.
Alignment touch up needed at all D-marked component!
|
Attachment 1: ETMY-ISCT_EISOL.jpg
|
|
9917
|
Tue May 6 17:58:44 2014 |
ericq | Update | LSC | farther into CM | I took a look at the MC OLTF and AO path TFs with the fast agilent analyzer.
I played with the relative gain of the EOM and PZT, but couldn't really change the MC OLTF shape much without making the PC Drive RMS angry.
However, it turns out we have plenty of phase headroom to up the MC UGF from ~100kHz to ~180, with about 40 degrees of phase margin and ~7dB of gain margin. As I write this, PC drive RMS is around 1.1, and FSS Fast at 5.6, so I think the extra gain is fine for now.
This pushes up and smoothens out the gain peaking in the AO path; see this figure:

(why does ELOG hate my python plots?! argggg)
Rana's rule of thumb was "We need at least +3dB MC loop gain at our CM servo UGF," so it looks like high tens of kHz bandwidth may be doable from the AO standpoint.
RXA: No, no, no, no, no, noooo. Rana said we need a gain of 3-10 at the CM UGF, not +3 dB. |
9916
|
Tue May 6 10:31:58 2014 |
jamie | Update | CDS | c1ioo dolphin fiber |
Quote: |
I put label at the dolphin fiber end at 1X2 today. After this I had to reset it, but it failed.
|
If by "fail" you're talking about the c1oaf model being off-line, I did that yesterday (see log 9910). That probably has nothing to do with whatever you did today, Steve. |
9915
|
Tue May 6 10:22:28 2014 |
steve | Update | CDS | c1ioo dolphin fiber |
Quote: |
Steve and I nicely routed the dolphin fiber from c1ioo in the 1X2 rack to the dolphin switch in the 1X4 rack. I shutdown c1ioo before removing the fiber, but still all the dolphin connected models crashed. After the fiber was run, I brought back c1ioo and restarted all wedged models. Everything is green again:

|
I put label at the dolphin fiber end at 1X2 today. After this I had to reset it, but it failed. |
Attachment 1: dolphin1x2.png
|
|
9914
|
Tue May 6 09:46:50 2014 |
steve | Update | PEM | cleaning day | Keven and Steve,
We cleaned around the Vertex chambers, PSL and MC2 floor areas.
|
Attachment 1: 120d.png
|
|
9913
|
Tue May 6 03:17:15 2014 |
rana | Update | LSC | farther into CM | Yes, we still need to do these things, day team. Please tune up the MC loop first, before anything else.
Quote: |
Next steps:
- Check CM board boosts turn on politely (Transients, TFs)
- Use fast spectrum analyzer to check MC loop gain out to a few MHz. (The bump in the tens of kHz should be fixed / moved higher)
- Think about noise performance of, say, REFLDC, ASDC, RF AS signals, etc. in the PRY case, figure out which one to use first.
- We may want to first focus on directly locking the arm on an RF signal, figure out gains etc. and then figure out how to do DC->RF handoff nicely, or if high bandwidth DC signal control is even feasible.
|
|
9912
|
Tue May 6 02:48:50 2014 |
Jenne | Update | LSC | AO path engaged with AS55 as error signal for Yarm locking | [Rana, Jenne]
This evening, we were able to lock the Yarm through the common mode board, using AS55 as our error signal. Our final UGF is about 5kHz, with 60 degrees of phase margin.
Before dinner, Rana switched the input of the CM board's REFL1 input to be AS55I rather than POY11Q, in the hopes that it would have better SNR. Demod phase of AS55 was measured to be 14 deg for optimum Yarm->I-phase and has been set to 0 degrees. Since the POY demod phase had been 90 degrees, which puts in a minus sign, and now we're using 0 deg which doesn't have a minus sign, we're using the plus (instead of minus) polarity of the CM board.
We re-allocated gains to help lower the overall noise by moving 15dB from the CM board AO gain slider to the MC IN2 gain slider, so we weren't attenuating signals.
We see, by taking loop measurements even before engaging the AO path (so, just the digital loop portion) that we've gained something like 20 degrees of phase margin! We think that about 5 degrees is some LSC loop re-shaping of the boost filter. We weren't sure why there was a hump of extra gain in the boost filter, so we've created a new (FM8) boost filter which is just a usual resonant gain: resgain(16.5,7,50)
The cm_down and cm_step scripts in ..../scripts/PRFPMI/ were modified to reflect the settings below, and their current states are included in the tarball attached.
Also, throughout our endeavors this evening, the PC fast rms has stayed nice and low, so we don't suspect any EOM saturation issues.
Now our Yarm digital servo has a gain of -0.0013, with FMs 2, 4, 5, 7, 8 engaged (2, 7, 8 are triggered).
Our final CM board settings are:
REFL1 gain = +22dB
offset = -2.898V
Boost = enable
Super Boost = 0
option = disable
1.6k:79 coupled cavity compensator = enabled
polarity = plus
option = disable
AO gain = 15dB
limiter = enable
MC board: IN1 gain = 18dB, IN2 gain = 0dB.
Here is a measurement of the Common Mode MCL/AO crossover. The purple/orange trace here is after/before the boost was engaged.

We also have a measurement of the total loop gain, measured with the SR785. The parameter file, as well as the python script to get the data, are in the tarball attached. Noteably, the excitation amplitude was 500mV, whereas Q and Rana yesterday were using 5 or 8 mV. We aren't sure why the big change was necessary to get a reasonable measurement out. This measurement is with the boost enabled.

Finally, here is a measurement of the MC error point spectra, with the CM boost on, after we reallocated the gains. There's a giant bump at several tens of kHz. We need to actually go out with the fast analyzer and tune up the MC loop.

|
Attachment 2: zipped.tgz
|
|